20200617 SX On-Sky Checkout. Can we operate with the gap contamination at act 358?

People on Zoom

- David, John H, Dave T. Greg, Guido, Julian, Shane, Doug, Joe, Xianyu

Opening

02:59 Chamber open - Wind at about 19 m/s

03:07 Going to pointing star

03:25 OSCO preset - OK till optimize gain failed

04:06 Loop bin1 closed. - optimize gain fail because failed to load distrubance (Xianyu) - manually set gains to 1.1, 0.55, 0.55 - TN, 040611, 040911 - Luci image 0001 - 1 to 1.2 seeing

04:36 Closed loop with LUCI rotator at 340 degrees - TN 043905, 043950 - LUCI 0002 - 1 - 1.2" seeing

05:02 ESM preset to M92 - script says 9.5 mag, measure 12.9. Pick wrong Ref Star - manually set gain to 2.0. 1.0, 1.0 - TN 050925, - Dave have to manaully put N375 camera in LUCI1 due to script problem. - Collect H data

05:30 Preset to M92 with 15.5 mag Ref Star - 16.6 mag measured ref star - Gain 1.2, 0.7, 0.7 - loop opened. Had to recenter on pyrmid via the PSF GUI and the reclose. - data collection continuing - loop opened - gains 0.8. 0.5, 0.5

06:10 OSCO -

-- DougMiller - 17 Jun 2020
Topic revision: r1 - 17 Jun 2020, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback